OpenStudy (anonymous):

Sreenivas decides to buy a nice horse. He pays $600 for it, and he is very content withthis strong animal. After a year, the value of the horse has increased to $700 and hedecides to sell the horse. But, already a few days later he regrets his decision to sellthe beautiful horse, and he buys it again. Unfortunately, he has to pay $800 to get itback, so he loses $100. After another year of owning the horse, he finally decides tosell the horse for $900. What is the overall profit or lose the man makes ?

ganeshie8 (ganeshie8):

money going out : -ve money coming in : +ve overall profit = -600 + 700 - 800 + 900 = 200

OpenStudy (anonymous):

700-600+900-800=200

OpenStudy (anonymous):

the two sales are independent of each other it doesn't matter whether he bought and sold one horse or two